What is the exact value of the sec 0 if (5,2) is on the terminal side of 0 in a standard position?
Answer by stanbon(75887)

What is the exact value of the sec(x) if (5,2) is on the terminal side of
angle "x" in a standard position?
----------------------------
If x = 5 and y = 2, then r = sqrt(5^2+2^2) = sqrt(29)
-------------------
Since sec(x) = r/x, sec(x) = sqrt(29)/5
=============================================
Cheers,
Stan H.